Expand the following functiona using logarithmic properties. Some help with B and C would be great! Thanks!

(6) f(x) = ln ( 201)), assume 27 –1 > 0. (c) f(x) = In 3x5e+2+2) ), assume 2 > 0 (2x + 1)2
Answer #1

6 21-100 fou = ln ets (2x-1) x2 so en (ex-5 ( 2x - 2) 3) - ln (1²) for Antar . = ln(a)-in (byt en cet-s) + ln (28-13 - en (x²